OpenStudy (anonymous):

Solve the equation x2 = 9. A. 3 B. 4.5 C. 18 D. 81

OpenStudy (rea201):

I believe you mean x^2=9. In that case the correct answer is A. x=3 because 3*3=9

OpenStudy (anonymous):

thank u so much!!!!

OpenStudy (rea201):

No problem

OpenStudy (anonymous):

if thats not what you mean then the answer is 4.5

OpenStudy (anonymous):

4.5 * 2 = 9